OpenStudy (jjenk):

Is x - 8 a factor of the function f(x) = -2x3 + 17x2 - 64? Explain.

OpenStudy (cwrw238):

if x - 8 is a factor then f(8) = 0 by the Factor Theorem f(8) = -2(8)^3 + 17*(8)^2 - 64 = 0 so x - 8 is a factor